A stock is expected to pay a dividend of $2.75 the end of the year (that is, D1 = $2.75), and it should continue to grow at a constant rate of 3% a year. If its required return is 15%, what is the stock's expected price 3 years from today? Round your answer to two decimal places.
Answer #1

stock price today is

=D1/(r-g)

=2.75/(15%-3%)

=22.9167

the stock's expected price 3 years from today

=stock price today*(1+g)^n

=22.9167*(1+3%)^3

=25.04

the above is answer..